# Battery Capacity Knee-Onset Identification and Early Prediction Using Degradation Curvature

### v2

@@ 3.2 Knee and knee-onset identification
As suggested by yeh2016matrix , it is empirically shown that given the matrix profile and the matrix profile index, the resulting corrected arc curve contains information about a possible regime change at each location of a time series. Therefore, we employ a similar approach to knee and keen-onset identification. Specifically, our goal is to obtain three time series, i.e., matrix profile, matrix profile index, and corrected arc curve, to represent a time-series approximated curvature that is produced in the previous section.
@@ 3.3.1 Toyota research institute dataset
The first battery dataset used in this work was generated by Toyota Research Institute in collaboration with Stanford University and Massachusetts Institute of Technology severson2019data attia2020closed , in which there are 169 lithium iron ferrous phosphate (LFP)/graphite cylindrical cells manufactured by A124 Systems (model APR18650M1A, 1.1 Ah nominal capacity). The total 169 cells are from 4 batches (”2017-05-12”, ”2017-06-30”, ”2018-04-12”, and ”2019-01-24”) with batch name denoting the date when the experiment was started. The cells were charged with a one-step or multi-step fast-charging protocol from 0% to 80% SoC and then charged with a uniform 1C constant current–constant voltage (CC-CV) charging step from 80% to 100% SoC. Subsequently, cells were discharged identically at 4C rate to 0% SoC. All cells were tested in an environmental chamber at a constant ambient temperature of 30 ${}^{\circ}$. The cells were cycled until they reached the end of life (EoL) threshold, i.e., 80% of their initial nominal capacity. Time-series cell voltage, current, and (surface) temperature in each cycle were continuously measured, while two battery health metrics, i.e., rated capacity (4C discharge, 30 ${}^{\circ}$) and internal resistance (±$\pm 3.6$ C pulse current, 30 or 33 ms pulse width, 80% SoC) were measured per cycle. All the cells in this dataset have knee occurrence on their capacity fade curves before the end of the experiment.
@@
The knee occurrence in this dataset is caused by lithium plating due to loss of delithiated negative electrode active material. Specifically, at high rates of loss of delithiated negative electrode active material (LAM), the negative electrode capacity eventually falls below the remaining lithium inventory in a cell. Then, the negative electrode will not be able to accommodate all the lithium from the positive electrode during charge, which leads to irreversible lithium plating and a resulting loss of lithium inventory (LLI). The loss of delithiated negative electrode active material, together with LLI, contribute to accelerated capacity fade, at which a knee also occurs. Moreover, a higher charge C-rate also accelerates the occurrence of the knee in this dataset. Lithium plating-induced knees are commonly observed or hypothesized in commercial LFP/graphite cells lewerenz2017post ansean2017operando lewerenz2017differential severson2019data attia2020closed .
@@ 3.3.2 Sandia national lab dataset
The second battery dataset used in this work was generated by Sandia National Laboratories preger2020degradation , in which there are 32 lithium nickel manganese cobalt oxide (NMC)/graphite cylindrical cells from LG Chem (model 18650HG2, 3Ah). The NMC cells were cycled at three different ambient temperatures (15 ${}^{\circ}$, 25 ${}^{\circ}$, and 35 ${}^{\circ}$) with different depths of discharges (0-100%, 20-80%, and 40-60%) and discharge currents (0.5C, 1C, 2C, and 3C). All the NMC cells were identically charged at 0.5C rate. To reduce the effect of manufacturing variances, at least 2 cells were tested for each combination of ambient temperature, depth of discharge, and discharge current (12 combinations). The cells were cycled beyond the end of life, defined as when they reach 80% of their initial nominal capacity. Time-series cell voltage, current, and (surface) temperature in each cycle were continuously measured, while one battery health metric, i.e., rated capacity (0.5C discharge, the same ambient temperature as that in each cycling test) was measured periodically. Note that 4 cells that were cycled with 40-60% depth-of-discharge are excluded from this work due to the fact that neither the knee nor EoL (i.e., 80% of their initial nominal capacity) was experienced before the end of the experiment; 5 cells that were cycled with 20-80% depth-of-discharge and one cell that was cycled with 0-100% depth-of-discharge are also excluded from this work due to the fact that their discharge capacity data is highly corrupted. In the end, 22 cells that have knee occurrence on their capacity fade curves before the end of the experiment are included in this work.
@@
The knee occurrence in this dataset is attributed to resistance growth that is caused by the growth of side reaction products (e.g., solid electrolyte interphase (SEI)) on the surface of the electrode preger2020degradation . The NMC/graphite cells in this dataset have discharge voltage-capacity curves that are relatively flat at high voltage levels and relatively steep at low voltage levels. At the beginning of life, the discharge ends within the steep region of the voltage-capacity curve. However, as the overpotential increases due to resistance growth during aging, the voltage-capacity curve is pushed downwards. As a result, a cell reaches its lower cut-off voltage more quickly, and the discharge will eventually end within the flat region of the voltage-capacity curve, making the discharge capacity highly sensitive to increasing resistance. As a result, the discharge capacity fade accelerates, assuming a linear resistance growth rate, which leads to a capacity knee. The resistance-induced knees are commonly observed or hypothesized in cells made of oxide-based cathode materials, such as NMC, as they often operate above the stability window of the electrolyte preger2020degradation ma2019hindering .
@@ 3.3.3 Synthetic dataset
The knee occurrence can also be caused by particle cracking attia2022knees . Specifically, the intercalation and deintercalation of lithium during cycling can induce alternating mechanical stress within the electrodes, which results in particle cracking. As the cracks propagate, new surfaces are created for SEI growth, which accelerates LLI. The accelerated LLI contributes to accelerated capacity fade, at which a knee occurs. The particle cracking-induced knees have been observed or hypothesized in cells with cathodes made of NMC pfrang2018long and NCA willenberg2020high .
@@
In order to verify the effectiveness of our proposed method in identifying capacity knees caused by other degradation mechanisms and their interactions, synthetic battery degradation data is generated using physics-based models. Specifically, underlying battery states (e.g., lithium concentrations) are simulated using a Doyle-Fuller-Newman (DFN) model doyle1995use . To produce particle cracking-induced knees, two degradation mechanisms (i.e., SEI growth and particle cracking) at the negative electrode are coupled with the DFN model in Python Battery Mathematical Modeling (PyBaMM) library sulzer2021python .
@@
The DFN model parameters (i.e., electrode parameters, electrolyte parameters) are taken from Chen et al. chen2020development for a cylindrical LGM50 cell that is comprised of an NMC 811 positive electrode and a graphite-$\text{SiO}_{\text{x}}$ negative electrode. The values of electrode and electrolyte parameters are listed in the supplementary information of Ref. o2022lithium . The parameters of the degradation models in PyBaMM are taken from multiple sources and listed in the supplementary information of Ref. o2022lithium . In particular, the cracking rate in Paris’law is varied in order to generate synthetic capacity knee for verification purposes. Note that the battery degradation process was simulated with 30 times the standard particle cracking rate (i.e., $3.9\times 10^{-20}$ purewal2014degradation ).
@@
The cylindrical LGM50 cell used in this work has a nominal capacity of 5 Ah with a lower voltage cut-off of 2.5 V and an upper voltage cut-off of 4.2 V. The ambient temperature is assumed to be constant at 25 ℃. The cell is charged with a 1C constant current-constant voltage (CC-CV) charging to 4.2 V and a current cut-off of C/100 (50 mA) followed by a rest for 5 min. The cell is subsequently discharged at 1C to 2.5 V with a current cut-off of C/100 (50 mA) and then at rest for 5 min. The simulation is set to terminate if either 1200 cycles or 80% of the initial nominal capacity is reached. The discharge capacity is measured by integrating discharge current over time from 100% SoC to the cut-off voltage.
@@ 4 Results and discussion
4.1 Results of knee and knee-onset identification
As a fundamental step prior to addressing capacity knee-related battery problems, one would like to accurately identify knee-onset and knee on the capacity fade curve and then investigate the empirical relationship between knee-onset and end of life (80% of initial nominal capacity), and between knee and end of life. In this work, we first validate our proposed capacity knee identification method on the two experimental degradation datasets of different battery chemistry (LFP and NMC) and then on the synthetic degradation dataset. In addition, we also benchmark our capacity knee identification method to the state-of-the-art knee identification method in the literature, i.e., the double Bacon-Watts model fermin2020identification .
4.1.1 Toyota research institute dataset
With in total 169 LFP cells in the Toyota Research Institute dataset, capacity knee-onset and capacity knee were identified for each cell using the proposed capacity knee identification method. For these, we found strong linear correlations between both knee-onset and end of life ($\rho=1.0$), and between knee and end of life ($\rho=0.992$), as shown in Fig. 3. In Table 3 and 4, we also compare the knee and knee-onset identification performance of our proposed method with identification using the double Bacon-Watts model. The performance improvement in percentage is calculated with the state-of-the-art method as benchmark, and shows that our proposed method outperforms the state-of-the-art method measured by Pearson’s correlation coefficient. Specifically, there is a 0.6% performance improvement in knee identification and a 1.5% performance improvement in knee-onset identification on the Toyota Research Institute dataset.
Apart from the strong correlations, by referring to the diagonal line [solid black line] in Fig. 3, it can be concluded that almost all the cells have knees occurring before the end of life. Moreover, the capacity knee method identified both earlier knees and earlier knee-onsets than those identified using the double Bacon-Watts model.
@@ 4.1.1 Toyota research institute dataset
Table 3: Battery capacity knee identification performance
Method Knee identification performance measured by Pearson’s r Performance improvement (%) Toyota Research Institute Sandia National Lab Toyota Research Institute Sandia National Lab Double Bacon-Watts model 0.994 0.125 / / Proposed method 1.000 0.710 0.6% 468%
Table 4: Battery capacity knee-onset identification performance
Method Knee-onset identification performance measured by Pearson’s r Performance improvement (%) Toyota Research Institute Sandia National Lab Toyota Research Institute Sandia National Lab Double Bacon-Watts model 0.977 0.180 / / Proposed method 0.992 0.712 1.5% 296%
4.1.2 Sandia national lab dataset
With the additional 22 NMC cells in the Sandia National Lab dataset, we again found strong linear correlations between knee-onset and end of life ($\rho=0.712$), and between knee and end of life ($\rho=0.71$) using the proposed capacity knee identification method, as shown in Fig. 4. However, neither the knee-onset nor the knee identified using the double Bacon-Watts model shows strong correlations with the end of life. As shown in Table 3 and 4, the knee and knee-onset identification performance of our proposed method again outperforms that of the state-of-the-art (i.e., the double Bacon-Watts model). Specifically, the Pearson correlation coefficient in knee identification is 468%, and in knee-onset identification it is 296%.
In order to find out the possible cause of the weak correlations between both knee-onsets and knees, identified using the double Bacon-Watts model and end of life, we compared the knee-onsets and the knees obtained using the double Bacon-Watts model and the proposed capacity knee identification method, respectively for a sample NMC cell [No.10] in the Sandia National Lab dataset. As illustrated in Fig. 5b, the sample cell exhibits convex capacity fade in the first degradation phase (i.e., from the beginning of life to the knee-onset point) instead of relatively linear capacity fade in Fig. 2b. The double Bacon-Watts model failed to identify both knee and knee-onset [green lines], while our proposed method successfully identified both knee and knee-onset [blue lines] on this NMC cell. The reason for the failure of the double Bacon-Watts model is that by setting the initial values of the model parameters in Table 1, assumptions regarding the capacity fade curve hold in the Toyota Research Institute dataset but not in the Sandia National Lab dataset. In contrast, our proposed curvature-based identification method does not require any assumptions of the capacity fade curve, instead, it identifies knee and knee-onset points, given approximated degradation curvature. As a matter of fact, almost all the NMC cells from the Sandia National Lab dataset exhibit convex capacity fade instead of linear capacity fade in the first degradation phase, unlike the LFP cells in the Toyota Research Institute dataset.
Furthermore, by referring to the diagonal line [solid black line] in Fig. 4, it can be seen that there are NMC cells in the Sandia National Lab dataset that have knees occurring both before and after the end of life, which motivates the need for classifying retired batteries based on whether or not the knee onset and the knee itself has occurred in their first lives.
Figure 4: The relationship between knee-onset and end of life (left), and knee and end of life (right) for the Sandia National Lab dataset.
@@ 4.1.2 Sandia national lab dataset
4.1.3 Synthetic dataset
In order to verify the effectiveness of our proposed method to identify a capacity knee caused by particle cracking, the synthetic battery degradation data generated using physics-based models is used. As a measure of the rate of change of degradation rate, the approximated curvature together with its corresponding CAC with respect to cycle number for the simulated LGM50 cell is plotted in Fig. 6a. A significant fluctuation of the approximated curvature can again be observed in the second state of which the boundaries [blue lines] were successfully inferred by the proposed identification method. With a state change at each boundary, very few arcs cross over the boundary as most of the approximated curvature subsequences should find their nearest neighbors within the same state. Therefore, the height of the CAC should be the lowest at each boundary where the LGM50 cell degradation process changes from one state to another, as shown at the bottom of Fig. 6a. The measured discharge capacity fade curves are shown in Fig. 6b. It can be seen that the knee is observable after approximately 680 cycles, and is followed by a sudden failure at 980 cycles due to the fact that the porosity at the negative electrode-separator interface reached zero.

### v3

@@ 3.2 Knee and knee-onset identification
As suggested by [18], it is empirically shown that given the matrix profile and the matrix profile index, the resulting corrected arc curve contains information about a possible regime change at each location of a time series. Therefore, we employ a similar approach to knee and keen-onset identification. Specifically, our goal is to obtain three time series, i.e., matrix profile, matrix profile index, and corrected arc curve, to represent a time-series approximated curvature that is produced in the previous section.
@@ 3.3.1 Toyota research institute dataset
The first battery dataset used in this work was generated by Toyota Research Institute in collaboration with Stanford University and Massachusetts Institute of Technology [23] [27], in which there are 169 lithium iron ferrous phosphate (LFP)/graphite cylindrical cells manufactured by A124 Systems (model APR18650M1A, 1.1 Ah nominal capacity). The total 169 cells are from 4 batches (”2017-05-12”, ”2017-06-30”, ”2018-04-12”, and ”2019-01-24”) with batch name denoting the date when the experiment was started. The cells were charged with a one-step or multi-step fast-charging protocol from 0% to 80% SoC and then charged with a uniform 1C constant current–constant voltage (CC-CV) charging step from 80% to 100% SoC. Subsequently, cells were discharged identically at 4C rate to 0% SoC. All cells were tested in an environmental chamber at a constant ambient temperature of 30C. The cells were cycled until they reached the end of life (EoL) threshold, set to 80% of their initial nominal capacity. Time-series cell voltage, current, and (surface) temperature in each cycle were continuously measured, while two battery health metrics, i.e., rated capacity (4C discharge, 30C) and internal resistance (±$\pm 3.6$ C pulse current, 30 or 33 ms pulse width, 80% SoC) were measured per cycle. All the cells in this dataset have knee occurrence on their capacity fade curves before the end of the experiment.
@@
The knee occurrence in this dataset is caused by lithium plating due to loss of delithiated negative electrode active material. Specifically, at high rates of loss of delithiated negative electrode active material (LAM), the negative electrode capacity eventually falls below the remaining lithium inventory in a cell. Then, the negative electrode will not be able to accommodate all the lithium from the positive electrode during charge, which leads to irreversible lithium plating and a resulting loss of lithium inventory (LLI). The loss of delithiated negative electrode active material, together with LLI, contribute to accelerated capacity fade, at which a knee also occurs. Moreover, a higher charge C-rate also accelerates the occurrence of the knee in this dataset. Lithium plating-induced knees are commonly observed or hypothesized in commercial LFP/graphite cells [28] [29] [30] [23] [27].
@@ 3.3.2 Sandia national lab dataset
The second battery dataset used in this work was generated by Sandia National Laboratories [31], in which there are 32 lithium nickel manganese cobalt oxide (NMC)/graphite cylindrical cells from LG Chem (model 18650HG2, 3Ah). The NMC cells were cycled at three different ambient temperatures (15C, 25C, and 35C) with different depths of discharges (0-100%, 20-80%, and 40-60%) and discharge currents (0.5C, 1C, 2C, and 3C). All the NMC cells were identically charged at 0.5C rate. To reduce the effect of manufacturing variances, at least 2 cells were tested for each combination of ambient temperature, depth of discharge, and discharge current (12 combinations). The cells were cycled beyond the end of life, defined as when they reach 80% of their initial nominal capacity. Time-series cell voltage, current, and (surface) temperature in each cycle were continuously measured, while one battery health metric, i.e., rated capacity (0.5C discharge at the same ambient temperature as that in each cycling test) was measured periodically. Note that 4 cells that were cycled with 40-60% depth-of-discharge are excluded from this work due to the fact that neither the knee nor EoL (i.e., 80% of their initial nominal capacity) was experienced before the end of the experiment; 5 cells that were cycled with 20-80% depth-of-discharge and one cell that was cycled with 0-100% depth-of-discharge are also excluded from this work due to the fact that their discharge capacity data is highly corrupted. In the end, 22 cells that have knee occurrence on their capacity fade curves before the end of the experiment are included in this work.
@@
The knee occurrence in this dataset is attributed to resistance growth that is caused by the growth of side reaction products (e.g., solid electrolyte interphase (SEI)) on the surface of the electrode [31]. The NMC/graphite cells in this dataset have discharge voltage-capacity curves that are relatively flat at high voltage levels and relatively steep at low voltage levels. At the beginning of life, the discharge ends within the steep region of the voltage-capacity curve. However, as the overpotential increases due to resistance growth during aging, the voltage-capacity curve is pushed downwards. As a result, a cell reaches its lower cut-off voltage more quickly, and the discharge will eventually end within the flat region of the voltage-capacity curve, making the discharge capacity highly sensitive to increasing resistance. As a result, the discharge capacity fade accelerates, assuming a linear resistance growth rate, which leads to a capacity knee. Resistance-induced knees are commonly observed or hypothesized in cells made of oxide-based cathode materials, such as NMC, as they often operate above the stability window of the electrolyte [31] [32].
@@ 3.3.3 Synthetic dataset
The knee occurrence can also be caused by particle cracking [6]. Specifically, the intercalation and deintercalation of lithium during cycling can induce alternating mechanical stress within the electrodes, which results in particle cracking. As the cracks propagate, new surfaces are created for SEI growth, which accelerates LLI. The accelerated LLI contributes to an accelerated capacity fade and a consequent knee occurrence. Particle cracking-induced knees have been observed or hypothesized in cells with cathodes made of NMC [33] and NCA [34].
@@
In order to verify the effectiveness of our proposed method in identifying capacity knees caused by other degradation mechanisms and their interactions, synthetic battery degradation data is generated using physics-based models. Specifically, underlying battery states (e.g., lithium concentrations) are simulated using a Doyle-Fuller-Newman (DFN) model [35]. To produce particle cracking-induced knees, two degradation mechanisms (i.e., SEI growth and particle cracking) at the negative electrode are coupled with the DFN model in Python Battery Mathematical Modeling (PyBaMM) library [36].
@@
The DFN model parameters (i.e., electrode parameters, electrolyte parameters) are taken from Chen et al. [37] for a commercial NMC 811/graphite-$\text{SiO}_{\text{x}}$ cylindrical cell from LG Chem (INR21700 M50, 5 Ah). The values of electrode and electrolyte parameters are listed in the supplementary information of Ref. [38]. The parameters of the degradation models in PyBaMM are taken from multiple sources and are also listed in the supplementary information of Ref. [38]. Here, three distinct cells with their cracking rates in Paris’ law being 10, 30, and 50 times the standard particle cracking rate (i.e., $3.9\times 10^{-20}$ [39]) were used to generate synthetic capacity knees for validation purposes.
@@
The cells used in this work have a nominal capacity of 5 Ah with a lower voltage cut-off of 2.5 V and an upper voltage cut-off of 4.2 V. The ambient temperature is assumed to be constant at 25C. The cells are charged with a 1C constant current-constant voltage (CC-CV) charging to 4.2 V and a current cut-off of C/100 (50 mA) followed by a rest for 5 minutes. The cells are subsequently discharged at 1C to 2.5 V with a current cut-off of C/100 (50 mA) and then at rest for 5 minutes. The simulation is set to terminate if either 1200 cycles or 80% of the initial nominal capacity is reached. The discharge capacity is measured by integrating discharge current over time from 100% SoC to the cut-off voltage.
@@ 4 Results and discussion
As a fundamental step prior to addressing capacity knee-related battery problems, one would like to accurately identify knee-onset and knee on the capacity fade curve and then investigate the empirical relationship between knee-onset and end of life (80% of initial nominal capacity), and between knee and end of life. In this section, we will first validate our proposed capacity knee identification method on the two experimental degradation datasets of different battery chemistry (LFP and NMC) and then on the synthetic degradation dataset. In the validation process, we will also benchmark our capacity knee identification method to the state-of-the-art knee identification method in the literature, i.e., the double Bacon-Watts model [14] and the numerical results of knee and knee-onset identification will also be presented. Lastly, a case study will be provided to demonstrate one promising application of knee-onset identification using our proposed method on experimental data of commercial LFP batteries.
4.1 Validation of the proposed method
4.1.1 Validation on Toyota research institute dataset
With in total 169 LFP cells in the Toyota Research Institute (TRI) dataset, capacity knee-onset and capacity knee were identified for each cell using the proposed capacity knee identification method. For these, we found strong linear correlations between both knee-onset and end of life ($\rho=1.0$), and between knee and end of life ($\rho=0.992$), as shown in Fig. 3. In Table 3, we also compare the knee and knee-onset identification performance of our proposed method with identification using the double Bacon-Watts model. It can be concluded that the identification performance of our proposed method outperforms that of the state-of-the-art (i.e., the double Bacon-Watts model).
Apart from the strong correlations, by referring to the diagonal line [solid black line] in Fig. 3, it can be concluded that all the cells have knees occurring before the end of life. It can also be noted that the capacity knee method identified both earlier knees and earlier knee-onsets than those identified using the double Bacon-Watts model.
@@ 4.1.1 Validation on Toyota research institute dataset
4.1.2 Validation on Sandia national lab dataset
With the additional 22 NMC cells in the Sandia National Lab (SNL) dataset, we again found clear linear correlations between knee-onset and end of life ($\rho=0.712$), and between knee and end of life ($\rho=0.71$) using the proposed capacity knee identification method, as shown in Fig. 4. However, neither the knee-onset nor the knee identified using the double Bacon-Watts model shows any clear correlations with the end of life.
Figure 4: The relationship between knee-onset and end of life (left), and knee and end of life (right) for 22 cells in the Sandia National Lab dataset.
As shown in Table 3, the knee and knee-onset identification performance of our proposed method again outperforms that of the state-of-the-art (i.e., the double Bacon-Watts model).
Table 3: Battery capacity knee and knee-onset identification performance
Method Knee identification performance measured by Pearson’s r Knee-onset identification performance measured by Pearson’s r TRI SNL TRI SNL Double Bacon-Watts model 0.994 0.125 0.977 0.180 Proposed method 1.000 0.710 0.992 0.712
In order to find out the possible cause of the weak correlations between both knee-onsets and knees, identified using the double Bacon-Watts model and end of life, we compared the knee-onsets and the knees obtained using the double Bacon-Watts model and the proposed capacity knee identification method for a sample NMC cell [No.10] in the SNL dataset. As illustrated in Fig. 5b, the sample cell exhibits convex capacity fade in the first degradation phase (i.e., from the beginning of life to the knee-onset point) instead of relatively linear capacity fade in Fig. 2b. The double Bacon-Watts model failed to identify both knee and knee-onset [green lines], while our proposed method successfully identified both knee and knee-onset [blue lines] on this NMC cell.
@@ 4.1.2 Validation on Sandia national lab dataset
The reason for the weak correlations between both knee-onsets and knees, identified using the double Bacon-Watts model and end of life, is that by setting the initial values of the model parameters in Table 1, assumptions regarding the capacity fade curve hold for all the LFP cells in the TRI dataset but not for all the NMC cells in the SNL dataset. In fact, the NMC cells in the SNL dataset exhibit highly diverse capacity fade curves. It is therefore not possible to find one set of initial values in Table 1 for all the NMC cells in the SNL dataset. It might be possible to improve the knee-onset and knee identification performance using the double Bacon-Watts model in the SNL dataset by first classifying the NMC cells into several groups based on the similarity level of their capacity fade curves and then setting different initial values in Table 1 for each group of NMC cells. However, this will inevitably incur additional work. In contrast, our proposed curvature-based identification method does not require any assumptions of the capacity fade curve, instead, it identifies knee and knee-onset points, given an approximated degradation curvature.
Furthermore, by referring to the diagonal line [solid black line] in Fig. 4, it can be seen that there are NMC cells in the SNL dataset that have knees occurring both before and after the end of life, which motivates the need for classifying retired batteries based on whether or not the knee onset and the knee itself has occurred in their first lives.
4.1.3 Validation on synthetic dataset
In order to verify the effectiveness of our proposed method to identify a capacity knee caused by particle cracking, we apply the method to the synthetic battery degradation data generated for the three LGM50 cells with cracking rates as in Table 4. It can be seen from Table 4 that both knee-onset and knee identified using the proposed method monotonically decrease with increasing cracking rates (from Cell 1 to Cell 3), and we again found a strong linear correlation between knee-onset and knee ($\rho=1.0$) using the proposed knee identification method. Also here, the knee-onset and knee identified using the double Bacon-Watts model show only a weak correlation ($\rho=0.213$).
Table 4: Battery capacity knee and knee-onset identification performance
Method Knee (cycles) Knee-onset (cycles) Identification performance measured by Pearson’s r Cell 1 (10 $\times$) Cell 2 (30 $\times$) Cell 3 (50 $\times$) Cell 1 (10 $\times$) Cell 2 (30 $\times$) Cell 3 (50 $\times$) Double Bacon-Watts model 863 804 541 164 537 364 0.213 Proposed method 840 686 455 358 291 191 1.000 Three cells with their cracking rates in Paris’ law being 10, 30, and 50 times the standard particle cracking rate (i.e., $3.9\times 10^{-20}$ [39])
As a measure of the rate of change of degradation rate, the approximated curvature together with its corresponding CAC is plotted versus the cycle number for the synthetic LGM50 cell 2 in Fig. 6a. A significant fluctuation of the approximated curvature can again be observed in the second state of which the boundaries [blue lines] were successfully inferred by the proposed identification method. With a state change at each boundary, very few arcs cross over the boundary as most of the approximated curvature subsequences should find their nearest neighbors within the same state. Therefore, the height of the CAC should be the lowest at each boundary where the LGM50 cell degradation process changes from one state to another, as shown at the bottom of Fig. 6a. The measured discharge capacity fade curves are shown in Fig. 6b. It can be seen that the knee is observable after approximately 680 cycles, and is followed by a sudden failure at 980 cycles due to the fact that the porosity at the negative electrode-separator interface reached zero.